# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v4

@@ IV-D4 Difficulties
a) Disadvantages of sensors with different principles: traditional sensors provide simple, limited, and low-dimensional data, posing challenges for multi-modal learning. Vision-based sensors and electronic skins, although offering high accuracy, are cost-prohibitive. Moreover, vision-based sensors cannot provide temperature information. b) Difficulties in data acquisition: tactile datasets are rare and heterogeneous, lacking the standardized, extensive repositories found in fields like vision. Data collection is particularly challenging, especially when gathering both tactile and visual information simultaneously, despite some efforts in developing simplified collection devices. c) Difficulties in inconsistent standards: there are various sensors on the market with inconsistent standards and principles. Even with similar imaging patterns, the design and calibration of vision-based tactile sensors result in a significant domain gap. Unlike standardized formats for visual and auditory data, tactile data formats vary widely among sensor manufacturers, making large-scale learning from heterogeneous sensors difficult and limiting the usefulness of publicly available tactile datasets.available tactile datasets.
@@ V-A Embodied Question Answering
For EQA task, the agent needs to explore the environment from a first-person perspective to gather information necessary to answer the given questions. An agent with autonomous exploration and decision-making capabilities must not only consider which actions to take to explore the environment but also determine when to stop exploring to answer questions. Existing works focus on different types of questions, some of which are shown in Fig. 11. In this section, we first introduce the existing datasets and then discuss related methods.
@@ V-A1 Datasets
Conducting robot experiments in real environments is often constrained by scenarios and robot hardware. As virtual experimental platforms, simulators offer suitable environmental conditions for constructing embodied question answering datasets. Training and testing models on datasets created in simulators significantly reduce experimental costs and enhance the success rate of deploying models on real machines. We briefly introduces nine embodied question answering datasets, which are summarized in Table IX.
@@
OpenEQA[269] 2024 Active EQA, Episodic Memory EQA ScanNet, HM3D Habitat Manual open-ended 1,000+
@@
EQA v1[262] is the first dataset designed for embodied question answering. Built on synthetic 3D indoor scenes from the SUNCG dataset[97] within the House3D[272] simulator, EQA v1 comprises four types of questions: location, color, color_room, and preposition. It features over 5,000 questions distributed across more than 750 environments. The questions are constructed via functional program execution, using templates to select and combine basic operations.
@@
SQA3D[267] simplified protocol (QA only) while still preserving the function of benchmarking embodied scene understanding, enabling more complex, knowledge-intensive questions and a much larger scale of data collection. Specifically, SQA3D offers a dataset with about 6,800 unique situations, 20,400 descriptions, and 33,400 diverse reasoning questions for these situations based on ScanNet[275] scenes.
Unlike previous datasets that explicitly specify target objects in questions, K-EQA[268] features complex questions with logical clauses and knowledge-related phrases, requiring prior knowledge to answer. It is built in AI2Thor and includes four types of questions: existence, counting, enumeration, and comparison. Each entity is mapped to a knowledge base and a knowledge graph is further constructed. In this work, the templates provided in IQA and MT-EQA are extended to a set of grammars. After specifying objects and logical relationships, knowledge graphs, scene graphs, etc. are introduced to generate questions and compute the ground truth answer. The resulting K-EQA dataset consists of 60,000 questions across 6000 different environment setups.
OpenEQA[269] is the first open-vocabulary dataset for EQA, supporting both episodic memory and active exploration cases. The episodic memory EQA (EM-EQA) tasks involve an agent developing an understanding of the environment from its episodic memory to answer questions, similar to VideoNavQA. In active EQA (A-EQA) tasks, the agent answers questions by taking exploratory actions to gather necessary information. Using ScanNet[275] and HM3D[276], human annotators constructed over 1,600 high-quality questions from more than 180 real world environments in Habitat.
Utilizing GPT4-V, HM-EQA[270] is constructed in the Habitat simulator using the Habitat-Matterport 3D dataset. It includes 500 questions across 267 different scenes, which can be roughly categorized into identification, counting, existence, status, and location. For consistency, each question in the dataset has four multiple choices.
@@ V-A2 Methods
The EQA task was first proposed by Das et al.[262]. In their work, the agent consists of four main modules: vision, language, navigation, and answering. These modules are primarily constructed using traditional neural building blocks: Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs). These modules undergo training in two phases. Initially, the navigation and response modules are trained independently on automatically generated expert navigation demonstrations using imitation or supervised learning. Subsequently, in the second phase, the navigation architecture is fine-tuned using policy gradients. Some subsequent works[277, 278] retain modules like the question answering module proposed by Das et al.[262] and improve the model. Additionally, Wu et al.[278] proposed integrating the navigation and QA modules into a unified SGD training pipeline for joint training, thereby avoiding employing deep reinforcement learning to simultaneously train the separately trained navigation and question answering modules.
There are also some works that attempt to increase the complexity and completeness of question answering tasks. From the perspective of task singularity, several works[263, 279] expand the task to include multiple objectives and multi-agent, respectively, making it necessary for the model to store and integrate the information obtained by the agent’s exploration through methods such as feature extraction and scene reconstruction. Taking into account the interaction between the agent and the dynamic environment, Gordon et al.[265] introduce the Hierarchical Interactive Memory Network. Control alternates between the planner, responsible for task selection, and the low level controllers, which carry out task execution. During this process, an Egocentric Spatial GRU (esGRU) is utilized to store spatial memory, enabling the agent to navigate and provide answers. There is also a limitation in previous works where agents are unable to use external knowledge to answer complex questions and lack knowledge of the explored parts of the scene. To address this, Tan et al.[268] propose a framework that leverages the neural program synthesis method and the table converted from the knowledge and 3D scene graphs, allowing the action planner to access object-related information. Additionally, an approach based on Monte Carlo Tree Search (MCTS) is used to determine the next location for the agent to move to.
LLMs/VLMs Methods. In recent years, large language models (LLMs) and visual language models (VLMs) have made continuous progress and demonstrated outstanding capabilities across various fields. Consequently, researchers attempt to apply these models to solve embodied question answering tasks without any additional fine-tuning.
@@
Majumdar et al.[269] explore using LLMs and VLMs for episodic memory EQA (EM-EQA) task and Active EQA (A-EQA) task. For EM-EQA task, they consider Blind LLMs, Socratic LLMs with language descriptions of the episodic memory, Socratic LLMs with descriptions of the constructed scene graph, and VLMs processing multiple scene frames. The A-EQA task extends EM-EQA methods with frontier-based exploration (FBE)[280] for problem-independent environment exploration. Some other works[270, 281] also employ frontier-based exploration method to identify areas for subsequent exploration and to build semantic maps. They end the exploration early utilizing conformal prediction or image-text matching to avoid over-exploration. Patel et al.[282] emphasize the question answering aspect of the task. They leverage multiple LLM-based agents to explore the environment and enable them to independently answer questions with “yes” or “no” answers. These individual responses are utilized to train a Central Answer Model, responsible for aggregating the responses and generating robust answers.
@@ V-A3 Metrics
The model’s performance is usually assessed based on two aspects: navigation and question answering. In the navigation, many works adhere to the approach introduced by Das et al[262]. and utilize indicators like the distance to the target object upon completion of navigation ($d_{T}$), the change in distance to target from initial to final position ($d_{\varDelta}$) and the smallest distance to the target at any point in the episode ($d_{min}$) to evaluate the performance of the model. They are tested at 10, 30, or 50 actions away from the target. There are also works that measure it based on indicators such as trajectory length, intersection-over-union score for target object ($IoU$), etc. For question answering, the evaluation mainly involves mean rank ($MR$) of the ground-truth answer in the answer list and accuracy of the answers. Recently, Majumdar et al.[269] introduce the concept of an aggregate LLM-based correctness metric (LLM-Match) to evaluate the accuracy of open-vocabulary answers. Additionally, they assess efficiency by incorporating the normalized length of the agent’s path as a weight for the correctness metric.
@@ V-B Embodied Grasping
Embodied interaction, in addition to engaging in question-answering interactions with humans, can also involve performing operations based on human instructions, such as grasping and placing objects, thereby completing interactions among the robots, humans and objects. Embodied grasping requires comprehensive semantic understanding, scene perception, decision-making, and robust control planning. The embodied grasping methods integrate traditional robotic kinematic grasping with large models such as LLMs [283] and vision-language foundation models [16], which enables agents to perform grasping tasks under multi-sensory perceptions, including visual active perception, language understanding and reasoning. Figure 12 (b) illustrates an overview of human-agent-object interaction, where the agent accomplishes embodied grasping tasks.
@@ V-B4 End-to-End Approaches
CLIPORT [297] is a language-conditioned imitation learning agent that combines the vision-language pre-trained model CLIP with the Transporter Net to create an end-to-end dual-stream architecture for semantic understanding and grasp generation. It is trained using a large number of expert demonstration data collected from virtual environments, enabling the agent to perform semantically guided grasping. Based on the OCID dataset, CROG [290] proposes a vision-language-grasping dataset and introduces a competitive end-to-end baseline. It leverages CLIP’s visual foundation capabilities to learn grasp synthesis directly from image-text pairs. Reasoning Grasping [291] introduces the first reasoning grasping benchmark dataset based on the GraspNet-1 Billion dataset and proposes an end-to-end reasoning grasping model. The model integrates multimodal large language models (LLMs) with vision-based robotic grasping frameworks to generate grasps based on semantics and vision. SemGrasp [292] is a method for semantic-based grasp generation that incorporates semantic information into grasp representations to generate dexterous hand grasp postures. It introduces a discrete representation aligning grasp space with semantic space, enabling the generation of grasp postures according to language instructions. To facilitate training, a large-scale grasp-text alignment dataset CapGrasp is proposed.
@@ V-B5 Modular Approaches
Figure 13: Chronological overview of the embodied agents. Different colors show different paradigms of embodied agents. MLM means the Multimodal Language Model which directly perceive the world and control the embodiment, VLM means the Visual-Language Model with the outer policy models, LLM + VLM means the LLM-based agent that perceives the world utilizing the VLM, and LLM means the Large-Language Model with visual context and outer policy models.
Figure 14: The overall architecture of the embodied agent. It consists of the embodied multimodal foundation model, visual perception module, high-level task planning module, and low-level action planning module.
@@ VI-A Embodied Multimodal Foundation Model
Embodied agents are required to recognize their environment visually, understand instructions audibly, and comprehend their own state to enable complex interactions and operations. This demands a model that integrates multiple sensory modalities and natural language processing capabilities to enhance the agent’s understanding and decision-making by synthesizing diverse data types. Thus, the Embodied Multimodal Foundation Model is emerging. Google DeepMind initiated research in the Robotics Foundation Model field eight years ago, continually exploring ways to scale models and data more effectively. Their findings revealed that leveraging foundation models and large, diverse datasets is the optimal strategy. They developed a series of works based on the Robotic Transformer (RT) [13], offering substantial insights for future research on embodied agents.
@@
To address the generalization limitations of embodied models, Google collaborated with 33 leading academic institutions to create the comprehensive Open X-Embodiment dataset [306], integrating 22 diverse data types. Using this dataset, they trained the universal large model RT-X. The diverse, cross-entity training data enabled RT-1 and RT-2 to achieve superior performance, demonstrating better generalization and new functionalities compared to models trained on domain-specific data. This has also promoted the participation of more open-source VLMs in the robotics community, such as EmbodiedGPT [307] based on LLaVA and RoboFlamingo [308] based on Flamingo. Although Open X-Embodiment provides a vast array of datasets, constructing datasets remains a challenge given the rapid evolution of embodied robotic platforms. To address this issue, AutoRT [309] created a system for deploying robots in new environments to collect training data, leveraging LLMs to enhance learning capabilities through more comprehensive and diverse data.
@@
Additionally, transformer-based architectures face inefficiency problems, especially since embodied models require long contexts that include information from vision, language, and embodied states, as well as memory related to the currently executed tasks. For instance, RT-2, despite its strong performance, has an inference frequency of only 1-3Hz. Several efforts are being made to improve this, such as deploying models on the edge through quantization and distillation, and using the Mixture of Experts (MoE) architecture to utilize only a subset of parameters during inference, resulting in faster inference speeds compared to dense models with the same number of parameters. Moreover, SARA-RT [310] employs a novel model fine-tuning method called “up-training,” which converts quadratic computational complexity into simple linear complexity, significantly enhancing model efficiency. This makes RT faster and more streamlined, with the potential for large-scale adoption of RT technology. In addition to RT, the Mamba architecture is used for tasks requiring longer sequences. RoboMamba [311] employs an efficient fine-tuning strategy called Policy Head, which masters various operational skills with minimal fine-tuning parameters (0.1% of the model) and time (20 minutes). Furthermore, its inference speed is seven times faster than existing robotic MLMs.
@@
Due to RT being based on generative models, it demonstrates significant advantages in understanding abstract tasks and high-level task planning, but it has shortcomings in low-level action planning. One issue is that generative models cannot precisely generate control action parameters, and another is the gap between high-level task planning and low-level action planning, which prevents agents from effectively replanning. To address this, Google proposed RT-Trajectory [312], which can automatically add robot trajectories, providing low-level, practical visual cues for the model to learn robot control strategies and helping robots generalize better. These technologies enable robots to make decisions faster, better understand their environments, and more effectively guide themselves in completing tasks. Moreover, building on the RT-2 framework, the Robot Transformer with Action Hierarchies (RT-H) introduced a hierarchical action framework, linking high-level task descriptions with low-level robotic motions through intermediate linguistic actions [4]. This innovation improved cross-task data sharing and enhanced task performance by 15%. RT-H also exhibits greater adaptability and generalization in dynamic environments, excelling in complex multi-task scenarios. Furthermore, the emergent capabilities of VLA models are currently confined to high-level planning and affordance tasks related to VLMs. They cannot exhibit new skills in low-level physical interactions and are limited by the skill categories in their datasets. Physical actions often exhibit clumsiness, such as unstable gripping or inaccurate placement. Future efforts should aim to incorporate reinforcement learning into the training framework of large models to enhance generalization. This would enable VLA models to autonomously learn and optimize low-level physical interaction strategies in real-world environments, resulting in more dexterous and precise execution of various physical actions.

### v5

@@ IV-D4 Difficulties
a) Disadvantages of different sensor types: Traditional sensors deliver simple and low-dimensional data, challenging multi-modal learning. Vision-based sensors and electronic skins, while highly accurate, are expensive. b) Data acquisition challenges: Collecting data, especially both tactile and visual simultaneously, is difficult despite some progress in developing simplified collection devices. c) Inconsistent standards: Tactile sensors operate with inconsistent standards and principles, hindering large-scale learning and limiting the usefulness of public datasets. There is a need for standardized and extensive datasets.
@@ V-A Embodied Question Answering
For EQA task, the agent needs to explore the environment from a first-person perspective to gather information necessary to answer the given questions. An agent with autonomous exploration and decision-making capabilities must not only consider which actions to take to explore the environment but also determine when to stop exploring to answer questions. Existing works focus on different types of questions, some of which are shown in Fig. 11. In this section, we will introduce the existing datasets, discuss the related methods, describe the metrics used to evaluate model performance, and address the remaining limitations of this task.
@@ V-A1 Datasets
Conducting robot experiments in real environments is often constrained by scenarios and robot hardware. As virtual experimental platforms, simulators offer suitable environmental conditions for constructing embodied question answering datasets. Training and testing models on datasets created in simulators significantly reduce experimental costs and enhance the success rate of deploying models on real machines. We briefly introduce several embodied question answering datasets, which are summarized in Table IX.
@@
OpenEQA[265] 2024 Active EQA, Episodic Memory EQA ScanNet, HM3D Habitat Manual open-ended 1,600+
@@
EQA v1[258] is the first dataset designed for EQA. Built on synthetic 3D indoor scenes from the SUNCG dataset[94] within the House3D[268] simulator, EQA v1 comprises four types of questions: location, color, color_room, and preposition. It features over 5,000 questions distributed across more than 750 environments. The questions are constructed via functional program execution, using templates to select and combine basic operations.
@@
SQA3D[263] simplifies protocol (QA only) while still preserving the function of benchmarking embodied scene understanding, enabling more complex, knowledge-intensive questions and a much larger scale of data collection. Specifically, SQA3D offers a dataset with about 6,800 unique situations, 20,400 descriptions, and 33,400 diverse reasoning questions for these situations based on ScanNet[271] scenes.
Unlike previous datasets that explicitly specify target objects in questions, K-EQA[264] features complex questions with logical clauses and knowledge-related phrases, requiring prior knowledge to answer. It is built in AI2Thor and includes four types of questions: existence, counting, enumeration, and comparison. Each entity is mapped to a knowledge base and a knowledge graph is further constructed. In this work, the templates provided in IQA[261] and MT-EQA are extended to a set of grammars. After specifying objects and logical relationships, knowledge graphs, scene graphs, etc. are introduced to generate questions and compute the ground truth answer. The resulting K-EQA dataset consists of 60,000 questions across 6000 different environment setups.
OpenEQA[265] is the first open-vocabulary dataset for EQA, supporting both episodic memory and active exploration cases. The episodic memory EQA (EM-EQA) tasks involve an agent developing an understanding of the environment from its episodic memory to answer questions, similar to VideoNavQA. In active EQA (A-EQA) tasks, the agent answers questions by taking exploratory actions to gather necessary information. Using ScanNet and HM3D[272], human annotators constructed over 1,600 high-quality questions from more than 180 real world environments in Habitat.
Utilizing GPT4-V, HM-EQA[266] is constructed in the Habitat simulator using HM3D. It includes 500 questions across 267 different scenes, which can be roughly categorized into identification, counting, existence, status, and location. For consistency, each question has four multiple choices.
@@ V-A2 Methods
The EQA task was first proposed by Das et al.[258]. In their work, the agent consists of four main modules: vision, language, navigation, and answering. These modules are primarily constructed using traditional neural building blocks: Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs). They undergo training in two phases. Initially, the navigation and response modules are trained independently on automatically generated expert navigation demonstrations using imitation or supervised learning. Subsequently, in the second phase, the navigation architecture is fine-tuned using policy gradients. Some subsequent works[273, 274] retained modules like the question answering module proposed by Das et al.[258] and improved the model. Additionally, Wu et al.[274] proposed integrating the navigation and QA modules into a unified SGD training pipeline for joint training, thereby avoiding employing deep reinforcement learning to simultaneously train the separately trained navigation and question answering modules.
There are also some works that attempted to increase the complexity and completeness of question answering tasks. From the perspective of task singularity, several works[259, 275] expanded the task to include multiple objectives and multi-agent, respectively, making it necessary for the model to store and integrate the information obtained by the agent’s exploration through methods such as feature extraction and scene reconstruction. Taking into account the interaction between the agent and the dynamic environment, Gordon et al.[261] introduced the Hierarchical Interactive Memory Network. Control alternates between the planner, responsible for task selection, and the low level controllers, which carry out task execution. During this process, an Egocentric Spatial GRU (esGRU) is utilized to store spatial memory, enabling the agent to navigate and provide answers. There is also a limitation in previous works where agents are unable to use external knowledge to answer complex questions and lack knowledge of the explored parts of the scene. To address this, Tan et al.[264] proposed a framework that leverages the neural program synthesis method and the table converted from the knowledge and 3D scene graphs, allowing the action planner to access object-related information. Additionally, an approach based on Monte Carlo Tree Search (MCTS) is used to determine the next location for the agent to move to.
LLMs/VLMs Methods. In recent years, LLMs and VLMs have made continuous progress and demonstrated outstanding capabilities across various fields. Consequently, researchers attempt to apply these models to solve embodied question-answering tasks without any additional fine-tuning.
@@
Majumdar et al.[265] explored using LLMs and VLMs for episodic memory EQA (EM-EQA) task and Active EQA (A-EQA) task. For EM-EQA task, they considered Blind LLMs, Socratic LLMs with language descriptions of the episodic memory, Socratic LLMs with descriptions of the constructed scene graph, and VLMs processing multiple scene frames. The A-EQA task extended EM-EQA methods with frontier-based exploration (FBE)[276] for problem-independent environment exploration. Some other works[266, 277] also employed frontier-based exploration method to identify areas for subsequent exploration and to build semantic maps. They ended the exploration early utilizing conformal prediction or image-text matching to avoid over-exploration. Patel et al.[278] emphasized the question answering aspect of the task. They leveraged multiple LLM-based agents to explore the environment and enable them to independently answer questions with “yes” or “no” answers. These individual responses are utilized to train a Central Answer Model, responsible for aggregating the responses and generating robust answers.
@@ V-A3 Metrics
The performance is usually assessed based on two aspects: navigation and question answering. In the navigation, many works adhered to the approach introduced by Das et al.[258] and utilized indicators like the distance to the target object upon completion of navigation ($d_{T}$), the change in distance to target from initial to final position ($d_{\varDelta}$) and the smallest distance to the target at any point in the episode ($d_{min}$) to evaluate the performance of the model. They are tested at 10, 30, or 50 actions away from the target. There are also works that measured it based on indicators such as trajectory length, intersection-over-union score for target object ($IoU$), etc. For question answering, the evaluation mainly involves mean rank ($MR$) of the ground-truth answer in the answer list and accuracy of the answers. Recently, Majumdar et al.[265] introduced the concept of an aggregate LLM-based correctness metric (LLM-Match) to evaluate the accuracy of open-vocabulary answers. Additionally, they assessed efficiency by incorporating the normalized length of the agent’s path as a weight for the correctness metric.
@@ V-B Embodied Grasping
Embodied interaction, in addition to question-answering interactions with humans, also involves performing operations based on human instructions, such as grasping and placing objects, thereby completing interactions among the robots, humans and objects. Embodied grasping requires comprehensive semantic understanding, scene perception, decision-making, and robust control planning. The embodied grasping methods integrate traditional robotic kinematic grasping with large models such as LLMs [279] and vision-language foundation models [14], which enables agents to perform grasping tasks under multi-sensory perceptions, including visual active perception, language understanding and reasoning. Figure 12 (b) illustrates an overview of human-agent-object interaction, where the agent accomplishes embodied grasping tasks.
@@ V-B4 End-to-End Approaches
CLIPORT [293] is a language-conditioned imitation learning agent that combines the vision-language pre-trained model CLIP with the Transporter Net to create an end-to-end dual-stream architecture for semantic understanding and grasp generation. It is trained using a large number of expert demonstration data collected from virtual environments, enabling the agent to perform semantically guided grasping. Based on the OCID dataset, CROG [286] proposes a vision-language-grasping dataset and introduces a competitive end-to-end baseline. It leverages CLIP’s visual foundation capabilities to learn grasp synthesis directly from image-text pairs. Reasoning Grasping [287] introduces the first reasoning grasping benchmark dataset based on the GraspNet-1 Billion dataset and proposes an end-to-end reasoning grasping model. The model integrates multimodal LLMs with vision-based robotic grasping frameworks to generate grasps based on semantics and vision. SemGrasp [288] is a method for semantic-based grasp generation that incorporates semantic information into grasp representations to generate dexterous hand grasp postures. It introduces a discrete representation aligning grasp space with semantic space, enabling the generation of grasp postures according to language instructions. To facilitate training, a large-scale grasp-text alignment dataset CapGrasp is proposed.
@@ V-B5 Modular Approaches
Figure 13: Chronological overview of the embodied agents. Different colors show different paradigms. MLM means the Multimodal Language Model which directly perceive the world and control the embodiment, VLM means the Visual-Language Model with the outer policy models, LLM + VLM means the LLM-based agent that perceives the world utilizing the VLM, and LLM means the Large-Language Model with visual context and outer policy models.
Figure 14: The architecture of the embodied agent based on embodied multimodal foundation model, which consists of visual perception module, high-level task planning module, and low-level action planning module.
@@ VI-A Embodied Multimodal Foundation Model
Embodied agents are required to recognize their environment visually, understand instructions audibly, and comprehend their own state to enable complex interactions and operations. This demands a model that integrates multiple sensory modalities and natural language processing capabilities to enhance the agent’s understanding and decision-making by synthesizing diverse data types. Thus, the Embodied Multimodal Foundation Model is emerging. Recently, Google DeepMind found that leveraging foundation models and large, diverse datasets is the optimal strategy. They developed a series of works based on the Robotic Transformer (RT) [11], offering substantial insights for future research on embodied agents.
@@
To address the generalization limitations of embodied models, Google collaborated with 33 leading academic institutions to create the comprehensive Open X-Embodiment dataset [302], integrating 22 diverse data types. Using this dataset, they trained the universal large model RT-X. This has also promoted the participation of more open-source VLMs for robotics, such as EmbodiedGPT [303] based on LLaVA and RoboFlamingo [304] based on Flamingo. Although Open X-Embodiment provides a vast array of datasets, constructing datasets remains a challenge given the rapid evolution of embodied robotic platforms. To address this issue, AutoRT [305] created a system for deploying robots in new environments to collect training data, leveraging LLMs to enhance learning capabilities through more comprehensive and diverse data.
@@
Additionally, transformer-based architectures face inefficiency problems because embodied models require long contexts that include information from vision, language, and embodied states, as well as memory related to the currently executed tasks. For instance, RT-2, despite its strong performance, has an inference frequency of only 1-3Hz. Several efforts have been made, such as deploying models through quantization and distillation. Moreover, improving the model framework is another viable approach. SARA-RT [306] employs more efficient linear attention, while RoboMamba [307] utilizes the mamba architecture, which is more efficient for long-sequence tasks. This enables it to achieve inference speeds seven times faster than existing robotic MLMs.
@@
Generative-model-based RT excels in high-level task understanding and planning but has limitations in low-level action planning due to the generative model’s inability to precisely generate action parameters and the gap between high-level task planning and low-level action execution. To address this, Google introduced RT-Trajectory [308], which provides low-level visual cues for learning robot control strategies by automatically adding robot trajectories. Similarly, building on the RT-2 framework, the Robot Transformer with Action Hierarchies (RT-H) incorporates a hierarchical action framework, linking high-level task descriptions with low-level robot motions through intermediate linguistic actions [4]. Furthermore, VLA models exhibit emergent capabilities only in high-level planning and affordance tasks related to VLMs. They fail to demonstrate new skills in low-level physical interactions and are constrained by the skill categories in their datasets, resulting in clumsy actions. Future research should integrate reinforcement learning into the training framework of large models to improve generalization, enabling VLA models to autonomously learn and optimize low-level physical interaction strategies in real-world environments, thus executing various physical actions more dexterously and accurately.